# Downscaling Neural Network for Coastal Simulations

### v2

@@ 4.3 Feature extraction
where $\mathbf{Z}$ is the extracted feature, $W_{1}$ and $W_{2}$ are learnable 2D convolutional parameters, $\sigma$ is the sigmoid function and $GP$ is the global pooling operation. We stack multiple such RCAN blocks to learn deep feature representation.
@@ 4.4 Spatiotemporal attention
The key component of the DNNCS is the spatiotemporal attention. Our goal is to jointly super-resolve the coarse coastal images spatially and temporally. PDEs typically govern spatiotemporal dynamical systems. Let us first consider a simplified wave equation as
$\small\frac{\partial^{2}u}{\partial t^{2}}=c^{2}\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}+\frac{\partial^{2}u}{\partial z^{2}}\right)$ (3)
@@
where $r_{\{h,v,d\}}$ is the horizontal, vertical, and depth pixel shuffle operator. It fuses the $t_{n}$-th and $t_{n+1}$-th wave equation features into one spatiotemporal map. $H(\cdot)$, $V(\cdot)$ and $D(\cdot)$ are the corresponding spatial maps that fuse the temporal features along the three dimensions. $G(\cdot)$ is the MultiHead Attention (MHA) operation. The learnable parameters are shared across horizontal-, vertical- and depth-view computation, which can find the nonlocal correlations across space and time. As shown in Figure 3, we fuse the features at adjacent frames to calculate the correlations among different directions. In order to record the pixel positional information, we propose to use learnable positional codes p to record the relative feature positions and attach them to the corresponding feature maps for computation.
@@
[…]
where $Concat$ concatenates the subset of features into one, $W_{\{Q,K,V\}}$ are query, key and value matrices. $LN$ is the layer normalization process, $FFN$ is the feed-forward network with multiple convolution layers for the attention output, and $d$ is the dimension of the feature map, which is used for normalization.
@@ 4.6 Losses for coastal downscaling
$\small L_{mae}=\frac{1}{H\times W\times C\times T}\sum_{i}^{H}\sum_{j}^{W}\sum_{k}^{C}\sum_{t}^{T}|Y_{i,j,k}(t)-Y^{\prime}_{i,j,k}(t)|$ (6)
@@
$\small L_{lp}=\frac{1}{H\times W\times C\times T}\sum_{i}^{H}\sum_{j}^{W}\sum_{k}^{C}\sum_{t}^{T}\frac{\left\|Y_{i,j,k}(t)-Y^{\prime}_{i,j,k}(t)\right\|^{2}}{\left\|Y(t)\right\|^{2}}$ (7)
@@
$\small L_{diff}=\left\|\frac{dY(t)}{dx}-\frac{dY^{\prime}(t)}{dx}\right\|^{2}+\left\|\frac{dY(t)}{dy}-\frac{dY^{\prime}(t)}{dy}\right\|^{2}+\left\|\frac{dY(t)}{dz}-\frac{dY^{\prime}(t)}{dz}\right\|^{2}$ (8)
@@ 5.1 Implementation Details
To generate sufficient data samples for model training and analysis, we use UTBEST [1] to simulate the tidal circulation in two locations: Bahamas (Bight of Abaco) and Galveston Bay. The geometry of both computational domains, bathymetry (bottom topography), and the coarsest computational meshes are illustrated in Figure 4. Finer meshes were obtained by subdividing each triangle into four via edge bisection, whereas the bathymetry on finer meshes was obtained by linear interpolation from the coarser mesh nodes to preserve the consistency of problem setup.
@@
At the open sea boundary, both test problems use the tidal elevation consisting of five constituents prescribed by the function
$\hat{\xi}$ $(t)=0.075\cos\left(\frac{t}{25.82}+3.40\right)+0.095\cos\left(\frac{t}{23.94}+3.60\right)$ (9)
$+0.1\cos\left(\frac{t}{12.66}+5.93\right)+0.395\cos\left(\frac{t}{12.42}\right)+0.06\cos\left(\frac{t}{12.00}+0.75\right)$
with time $t$ in hours and use a quadratic bottom friction $\tau_{bf}=C_{f}|\mathbf{q}|/H^{2}$ with $C_{f}=0.009$ (Bahamas) and $C_{f}=0.004$ (Galveston). The Coriolis parameter for Bahamas and Galveston is set to $3.19\times 10^{-5}~{}\mathrm{s}^{-1}$ and $7.07\times 10^{-5}~{}\mathrm{s}^{-1}$, respectively (see [1] for a more detailed description of the test problems).
The simulations are run for 24 hours with the time step for the lowest approximation order (piecewise constant DG spaces) and the lowest mesh resolution set to 20s for the Bahamas and 12s for the Galveston dataset. Output files with the water height $\xi$ and depth-integrated horizontal velocities $U$ and $V$ are generated in each time step for all configurations. Increasing the mesh resolution or the polynomial discretization order requires smaller time steps (we halve the time step for each mesh refinement and each order increment) and results in more output files. To provide paired LR and HR training data, we consider three levels of downscaling scenarios: size of simulation data, polynomial order of mesh reconstruction and the temporal sampling. Table 1 summarizes the data used for training, with the number of mesh elements shown in “Resolution” and the number of time steps in the “Order” column. We use three approximation orders and resolutions to produce corresponding coarse and fine simulation data at different temporal resolutions. That is, given three approximation orders and three different resolutions, we have a total of $3\times 3$ simulation cases for each location. Then, we render all data of their three views ($U$, $V$, $\xi$ planes) into grayscale images. Finally, all three views are merged as RGB images as the training data. We train one model on both the Bahamas and Galveston datasets for our evaluation.
Table 1: Summary of the data used in coastal simulation for Bahamas and Galveston. The “Resolution” value represents the number of mesh elements. In the “Order” column, we list the number of time steps in the 24-hour simulation for the corresponding polynomial discretization order.
@@
Resolution Order 0 Order 1 Order 2 Resolution Order 0 Order 1 Order 2
1696 4320 8640 17280 3397 1800 3600 7200
6784 8640 17280 34560 13588 3600 7200 14400
27136 17280 34560 69120 54352 7200 14400 28800
We randomly split the data into training, validation, and testing in $6:2:2$ ratios without overlapping. The rendered images are of identical dimensions, approximately $900\times 500$ pixels. We also mask out the land area and only focus on the water-covered subdomain downscaling. In the training phase, we learn 4$\times$ spatial downscaling and 2$\times$ temporal downscaling across different orders. We randomly select two adjacent LR coastal images ($\mathbf{X}(t_{n}),\mathbf{X}(t_{n+1})$) from lower resolutions to predict three corresponding HR images $\mathbf{Y}(t_{2n}),\mathbf{Y}(t_{2n+1}),\mathbf{Y}(t_{2n+2})$ from next higher-resolution. Then we crop the image pairs as patches of size $64\times 64$. We randomly apply horizontal & vertical flips, rotations, and frame reversal for data augmentation.
@@
  • 2.
    SSIM [78] measures the structural similarity between ground truth and estimation
  • 3.
    GMSD [81] measures the gradient differences for perceptual quality assessment
  • 4.
    LPIPS [83] measures the perceptual similarity between ground truth and estimation
To measure the model complexity, we use FLOPs, memory, and the number of parameters to compare the computation efficiency.
@@ 5.2 Comparison of the spatiotemporal downscaling with state-of-the-art methods
Visually, we report the downscaling results of different methods in Figure 5 for the Bahamas dataset and in Figure 6 for the Galveston Bay. Each plane is normalized between [0, 255] and shown as a single-channel grayscale image, and we also stack normalized three planes $U,V,\xi$ in order as RGB images. In Figure 5, the last row shows the ground truth. Denoting the simulation output for the n-th time step by $t_{n}$, we show the results for $t_{15}$ in the first column and the residual maps of $U$-, $V$- and $\xi$-planes between predictions and ground truth in columns 2, 3 and 4, respectively. The last column illustrates the residual maps between $t_{15}$ and $t_{30}$ to visualize the temporal changes. We can observe that 1) VDSR, RDN, EDVR, and PhySR wrongly predict the $V$-plane simulations, ST-Bicubic, SwinIR, and VRT mispredict the $U$-plane simulations, ResShift fails on both U- and $\xi$-planes. We can better minimize the simulation differences simultaneously for $U$-, $V$- and $\xi$-planes. 2) From the last column, we can see that ours can better match with the ground truth changes, while others produce large errors. In Figure 6 of the downscaling results on Galveston Bay (here, the ground truth is shown in the last column), we use white arrows to highlight the significant differences. We can see that 1) globally, ResShift and VRT enlarge the errors caused by ST-Bicubic (the green areas around the boundaries at time $t_{61}$). EDVR produces the wrong pattern at time $t_{41}$. 2) From the enlarged areas, we can see that ours can better mimic the changes of the solution around the corners. Others, like VDSR and PhySR* magnify the zigzag patterns. ResShift and VRT oversmooth the solution which visually causes color shifts.
@@
We report the number of parameters, FLOPs, and runtime for model complexity in Table 3. The FLOPs are computed with the input of LR size 180 × 180 and ×4 upsampling settings. We can see that ours has the third-lowest model complexity regarding the number of parameters and runtime. In general, the model performance and complexity contradict each other. To better visualize the trade-off comparisons among different methods, we summarize the information in Table 2 and Table 3 and visualize them in Figure 7. Ours is located at the bottom left corner, which indicates that it achieves the best balance between model complexity and reconstruction quality.
@@ 5.3 Ablation studies and analysis
The key module of our proposed DNNCS is the spatiotemporal attention. To visually understand its ability to capture the pixel information for simulation, we use Local Attribution Maps (LAM) [35] to visualize the influences of neighborhood pixels on the region of interest (ROI) in Figure 9. The idea is to calculate the path-integrated gradient along a gradually changing path from the downscalied result to the LR input. In Figure 9, we show the contribution’s pixel attribute map and region in both $t_{n}$ and $t_{n+1}$ LR inputs. The red box region is calculated and enlarged for comparison. We can see that EDVR and ours can explore wider regions, which means that they can extract a wider range of correlated features for calculation. RCAN can also explore larger regions because of its channel attention computation, however, it fails to show the differences caused by the temporal changes. On the contrary, we can observe that ours shows attribution changes (the red color ranges become larger from $t_{n}$ to $t_{n+1}$ LR image) because ours can explore the temporal changes and reflect the attention score on the different LR images.
@@
Figure 12: Plot of frame-by-frame RMSE scores by using different methods. It compares RMSE across 165 frames for different methods, evaluating spatial and temporal consistency. Temporal RMSE curves for RDN, EDVR and PhySR show prominent oscillations and higher overall error, indicating instability and poor motion coherence across frames. In contrast, spatial RMSE remains relatively flat, suggesting more consistent intra-frame quality. Our method (in pink for temporal and red for spatial) shows rapid convergence within the first 20 frames and maintains the lowest and most stable RMSE afterward, especially in the temporal domain, demonstrating superior consistency over time. There are 8 lines. It is just that some lines are very close to each other. I added some description to the caption and main text.
Meanwhile, as introduced in Section III, the proposed spatiotemporal attention explores the horizontal(H)-, vertical(V)- and depth(D)-dimensional feature extraction hence, we conduct the experiments to compare the combination of three sub-attention modules and report the results in Figure 10. On top, we show the RMSE values of using attention operation at different dimensions. “No attention” means no attention is used in the network. “H+V+D+Pos” means our final model with the positional encoded spatiotemporal attention module. We can conclude that 1) combining horizontal, vertical, and depth attention achieves the best performance, and 2) from the feature maps, we can observe that the module can explicitly capture the dominant feature in corresponding directions.

### v3

@@ 4.3 Feature extraction
where $\mathbf{Z}$ is the extracted feature, $W_{1}$ and $W_{2}$ are learnable 2D convolutional parameters, $\sigma$ is the sigmoid function and $GP$ is the global pooling operation. We stack multiple RCAN blocks to learn deep feature representations.
@@ 4.4 Spatiotemporal attention
The key component of the DNNCS is the spatiotemporal attention. Our goal is to jointly super-resolve the coarse coastal images spatially and temporally. Because PDEs govern the spatiotemporal dynamics we aim to model, we draw inspiration from the wave equation:
$\frac{\partial^{2}u}{\partial t^{2}}=c^{2}\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}+\frac{\partial^{2}u}{\partial z^{2}}\right)$ (3)
@@
where $r_{\{h,v,d\}}$ are the horizontal, vertical, and depth pixel shuffle operators. It fuses the $t_{n}$-th and $t_{n+1}$-th wave equation features into one spatiotemporal map. $H(\cdot)$, $V(\cdot)$ and $D(\cdot)$ are the corresponding spatial maps that fuse the temporal features along the three dimensions. $G(\cdot)$ is the MultiHead Attention (MHA) operation. The learnable parameters are shared across horizontal-, vertical- and depth-view computation, which can find the nonlocal correlations across space and time. As shown in Figure 3, we fuse the features at adjacent frames to calculate the correlations among different directions. In order to record the pixel positional information, we propose to use learnable positional codes p to record the relative feature positions and attach them to the corresponding feature maps for computation.
@@
$\begin{matrix}\!\begin{aligned} &Y=FFN(MHA(X))+X\\ &MHA(X)=X+Concat(head_{1},head_{2},...,head_{n})W_{O}\\ &head_{i}=softmax\left(\frac{LN(W_{Q}(X))LN(W_{K}(X))}{\sqrt{d}}\right)W_{V}(X)\end{aligned}\end{matrix}$ (5)
where $Concat$ concatenates the subset of features into one, $W_{\{Q,K,V\}}$ are query, key and value matrices. $W_{O}$ is the weighting matrix for total $O$ heads. $LN$ is the layer normalization process, $FFN$ is the feed-forward network with multiple convolution layers for the attention output, and $d$ is the dimension of the feature map, which is used for normalization.
@@ 4.6 Losses for coastal downscaling
$L_{mae}=\frac{1}{H\times W\times C\times T}\sum_{i}^{H}\sum_{j}^{W}\sum_{k}^{C}\sum_{t}^{T}|Y_{i,j,k}(t)-Y^{\prime}_{i,j,k}(t)|$ (6)
@@
$L_{lp}=\frac{1}{H\times W\times C\times T}\sum_{i}^{H}\sum_{j}^{W}\sum_{k}^{C}\sum_{t}^{T}\frac{\left\|Y_{i,j,k}(t)-Y^{\prime}_{i,j,k}(t)\right\|^{2}}{\left\|Y(t)\right\|^{2}}$ (7)
@@
$L_{diff}=\left\|\frac{dY(t)}{dx}-\frac{dY^{\prime}(t)}{dx}\right\|^{2}+\left\|\frac{dY(t)}{dy}-\frac{dY^{\prime}(t)}{dy}\right\|^{2}+\left\|\frac{dY(t)}{dz}-\frac{dY^{\prime}(t)}{dz}\right\|^{2}$ (8)
@@ 5.1 Tidal scenarios
To generate sufficient data samples for model training and analysis, we use the SYCL implementation [20, 21] of the UTBEST [1] model to simulate the tidal circulation in two locations: Bahamas (Bight of Abaco) and Galveston Bay. The geometry of both computational domains, bathymetry (bottom topography), and the coarsest computational meshes are illustrated in Figure 4; all tidal setups in this section are based on the validated ADCIRC [66] datasets. Finer meshes were obtained by subdividing each triangle into four via edge bisection, whereas the bathymetry on finer meshes was obtained by linear interpolation from the coarser mesh nodes to preserve the consistency of the problem setup between different mesh resolutions.
@@
The Bahamas test problem setup is based on [91]; it uses the tidal elevation consisting of five constituents (O1, K1, N2, M2, S2) at the open sea boundary, a quadratic bottom friction $\tau_{bf}=C_{f}|\mathbf{q}|/H^{2}$ with $C_{f}=0.009$, and a constant Coriolis force with the coefficient set to $3.19\times 10^{-5}~\mathrm{s}^{-1}$. The Galveston test also uses a quadratic friction with $C_{f}=0.004$ and a constant Coriolis force with the coefficient $7.07\times 10^{-5}~\mathrm{s}^{-1}$. In the following sections, the open boundary uses exactly the same forcing as the Bahamas test case, while in Section 5.4 we generate tidal components M2, N2, O1, S2, K1, P1, Q1, K2 to see the effect of changing forcing on the prediction. See [1] for a more detailed description of the test problems.
The simulations are run for 24 hours with the time step for the lowest approximation order (piecewise constant DG spaces) and the lowest mesh resolution set to 20s for the Bahamas and 12s for the Galveston dataset. Output files with the water height $\xi$ and depth-integrated horizontal velocities $U$ and $V$ are generated in each time step for all configurations. Increasing the mesh resolution or the polynomial discretization order requires smaller time steps (we halve the time step for each mesh refinement and each order increment) and results in more output files. The discontinuous Galerkin scheme uses Local Lax-Friedrichs flux and explicit strong stability preserving (SSP) Runge-Kutta time stepping schemes [42] of orders one, two, and three, chosen corresponding to the order of the polynomial discretization in space.
Table 1 summarizes the data used for training, with the number of mesh elements shown in “Resolution” and the number of time steps in the “Order” column. Given three approximation orders and three different resolutions, we have a total of $3\times 3$ simulation cases for each location. Then, we render all data of their three views ($U$, $V$, $\xi$ planes) into grayscale images. Finally, all three views are merged as RGB images (this is purely a trick for efficient data representation and to leverage standard multi-channel convolutional architectures) as the training data. We train one model on both the Bahamas and Galveston datasets for our evaluation.
[…]
Table 1: Summary of the data used in coastal simulation for Bahamas and Galveston. The “Elem.” value represents the number of mesh elements, ’Min/Max’ column contains the minimum/maximum element size in meters. In the “Order” column, we list the number of time steps in the 24-hour simulation for the corresponding polynomial discretization order, which refers to the polynomial degree used to approximate each solution component per mesh cell.
@@
Elem. Min/Max Order 0 Order 1 Order 2 Elem. Min/Max Order 0 Order 1 Order 2
1696 640/2400 4320 8640 17280 3397 200/4400 1800 3600 7200
6784 320/1200 8640 17280 34560 13588 100/2200 3600 7200 14400
27136 160/600 17280 34560 69120 54352 50/1100 7200 14400 28800
[…]
@@
  • SSIM [90] (Structural SIMilarity) measures the structural similarity between ground truth and estimation
  • GMSD [96] (gradient Magnitude Similarity Deviation) measures the gradient differences for perceptual quality assessment
  • LPIPS [98] (Learned Perceptual Image Patch Similarity) measures the perceptual similarity between ground truth and estimation
To measure the model complexity, we use FLOPs (floating point operations), memory, and the number of parameters to compare the computation efficiency.
@@ 5.2 Comparison of the spatiotemporal downscaling with state-of-the-art methods
Visually, we report the downscaling results of different methods in Figure 5 for the Bahamas dataset and in Figure 6 for the Galveston Bay. Each plane is normalized between [0, 255] and shown as a single-channel grayscale image, and we also stack normalized three planes $U,V,\xi$ in order as RGB images.
That is, the velocity components $U$, $V$, and surface elevation $\xi$ are independently normalized for neural network training. The RGB and grayscale visualizations are therefore normalized representations used to compare prediction accuracy, rather than direct physical quantities. In the RGB visualizations, each color channel corresponds to one variable ($U$, $V$, or $\xi$). A red-, green-, or blue-dominated pixel indicates a relatively larger error (or value) in the corresponding variable after normalization. Similarly, in the grayscale plots, brighter pixels indicate larger normalized magnitudes or errors. These colors are intended to highlight spatial error patterns and relative performance, not to convey absolute physical values.
In Figure 5, the last row shows the ground truth. Denoting the simulation output for the n-th time step by $t_{n}$, we show the results for $t_{15}$ in the first column and the residual maps of $U$-, $V$- and $\xi$-planes between predictions and ground truth in columns 2, 3 and 4, respectively. The last column illustrates the residual maps between $t_{15}$ and $t_{30}$ to visualize the temporal changes. We can observe that 1) VDSR, RDN, EDVR, and PhySR wrongly predict the $V$-plane simulations, ST-Bicubic, SwinIR, and VRT mispredict the $U$-plane simulations, ResShift fails on both U- and $\xi$-planes. We can better minimize the simulation differences simultaneously for $U$-, $V$- and $\xi$-planes. 2) From the last column, we can see that our method can better match with the ground truth changes, while other methods produce large errors. In Figure 6 of the downscaling results on Galveston Bay (here, the ground truth is shown in the last column), we use white arrows to highlight the significant differences. We can see that 1) globally, ResShift and VRT enlarge the errors caused by ST-Bicubic (the green areas around the boundaries at time $t_{61}$). EDVR produces the wrong pattern at time $t_{41}$. 2) From the enlarged areas, we can see that our method can better mimic the changes of the solution around the corners. Others, like VDSR and PhySR* magnify the zigzag patterns. ResShift and VRT oversmooth the solution which visually causes color shifts.
@@
We report the number of parameters, FLOPs, and runtime for model complexity in Table 3. The FLOPs are computed with the input of LR size $180\times 180$ and $\times 4$ upsampling settings for all comparisons. We can see that ours has the third-lowest model complexity regarding the number of parameters and runtime. In general, the model performance and complexity contradict each other. To better visualize the trade-off comparisons among different methods, we summarize the information in Table 2 and Table 3 and visualize them in Figure 7. Ours is located at the bottom left corner, which indicates that it achieves the best balance between model complexity and reconstruction quality.
@@ 5.3 Ablation studies and analysis
The key module of our proposed DNNCS is the spatiotemporal attention. To visually understand its ability to capture the pixel information for simulation, we use Local Attribution Maps (LAM) [43] to visualize the influences of neighborhood pixels on the region of interest (ROI) in Figure 9. The idea is to calculate the path-integrated gradient along a gradually changing path from the downscaling result to the LR input. In Figure 9, we show the contribution’s pixel attribute map and region in both $t_{n}$ and $t_{n+1}$ LR inputs. The red box region is calculated and enlarged for comparison. We can see that EDVR and ours can explore wider regions, which means that they can extract a wider range of correlated features for calculation. RCAN can also explore larger regions because of its channel attention computation, however, it fails to show the differences caused by the temporal changes. On the contrary, we can observe that ours shows attribution changes (the red color ranges become larger from $t_{n}$ to $t_{n+1}$ LR image) because ours can explore the temporal changes and reflect the attention score on the different LR images.
@@
Figure 12: Plot of frame-by-frame RMSE scores by using different methods. It compares RMSE across 165 frames for different methods, evaluating spatial and temporal consistency. Temporal RMSE curves for RDN, EDVR and PhySR show prominent oscillations and higher overall error, indicating instability and poor motion coherence across frames. In contrast, spatial RMSE remains relatively flat, suggesting more consistent intra-frame quality. Our method (in pink for temporal and red for spatial) shows rapid convergence within the first 20 frames and maintains the lowest and most stable RMSE afterward, especially in the temporal domain, demonstrating superior consistency over time.
Meanwhile, as introduced in Section III, the proposed spatiotemporal attention explores the horizontal(H)-, vertical(V)- and depth(D)-dimensional feature extraction hence, we conduct the experiments to compare the combination of three sub-attention modules and report the results in Figure 10. On top, we show the RMSE values of using attention operation at different dimensions. “No attention” means no attention is used in the network. “H+V+D+Pos” means our final model with the positional encoded spatiotemporal attention module. We can conclude that 1) combining horizontal, vertical, and depth attention achieves the best performance, and 2) the feature maps reveal how the module focuses on the most important pixels in each direction. Specifically, by visualizing the attention maps, we can see that the model assigns higher weights to pixels corresponding to dominant features along horizontal, vertical, or depth directions.
@@
5.4 Downscaling adaptation to changing forcing data
We note that the Galveston data may also include other tidal forcings, such as wind and river inflow. To evaluate the adaptability of our model, we simulate an additional Galveston dataset with modified tidal forcings and directly apply the pretrained model to this new dataset without further fine-tuning. As shown in Figure 13, the model exhibits similar performance across the two Galveston simulations, demonstrating that the proposed approach is simulation-agnostic. Table 5 further reports the quantitative downscaling results for both simulations, where only minor differences are observed across all evaluation metrics.
Figure 13: Visual comparison (RGB values) of downscaling performance on Galveston datasets. Given the original simulation (version 1), and the modified tidal forcings (version 2), the proposed model provides consistent downscaling visual quality.
Metric Resolution version 1 version 2 Resolution version 1 version 2
[…]
MAE 0.002 0.003 0.016 0.014
SSIM 0.972 0.970 0.979 0.975
GMSD 0.043 0.045 0.079 0.079
LPIPS 0.108 0.109 0.079 0.077
Table 5: Quantitative comparison of two Galveston simulations: version 1 with original simulation settings and version 2 with modified tidal forcings. The metrics are calculated on the average performance on the testing dataset without finetuning.
5.5 Storm surge during Hurricane Ike
To illustrate the use of our proposed method to a realistic flood scenario, we generate data using the Discontinuous Galerkin Shallow Water Equation Model (DG-SWEM) [57, 93]. This numerical model numerically approximates the SWE in (1), where the right hand side has been augmented with tidal forces, winds, and air pressure gradients. In addition, DG-SWEM has an advanced wetting and drying scheme to track the critical wet-dry interface during inundation [17]. The selected event to simulate is Hurricane Ike (2008) [14, 49], which led to historic storm surge levels on the Texas coast.
Figure 14: Computational domain for Ike storm surge simulations, bathymetry cut off at 50m below and 20m above sea level according to the North Atlantic Vertical Datum of 1988 (in meters).
With DG-SWEM, we simulate storm surge during this hurricane using winds and air pressures from high-resolution reanalysis data as described in [49], and tidal forcings from the tidal constituents (M2, N2, O1, S2, K1, P1, Q1, K2) obtained from the TPXO9 model [28]. We run DG-SWEM using a Local Lax-Friedrichs flux, polynomial order one, and a three stage - second order explicit Runge Kutta time stepping scheme. The computational domain in shown in Figure 14 and covers the entire ocean to the west of the 60 meridian. The simulations cover the time period of September 5, 2008 - September 14, 2008 (noon to noon), with a time step size of $0.5$ seconds. Two meshes were used for simulation – see Table 6 for details of the mesh resolutions and computational resources used in the simulation on the Frontera cluster at the Texas Advanced Computing Center (TACC). The two meshes and corresponding simulations differ only in mesh resolution, see [24] for further detail and validation of these datasets.
Table 6: Summary of the data used in Ike storm surge simulations. The “Elem.” value represents the number of mesh elements, ‘Min/Max’ column contains the minimum/maximum element size in meters. In the “# CPU cores“ column, we list the total number of Intel Xeon Platinum 8280 ‘Cascade Lake’ CPU cores (56 cores per node of Frontera cluster) used in the simulation and ‘Wallclock time’ is the corresponding total wallclock time of the simulation.
Elem. Min/Max (meters) # CPU cores Wallclock time
3,832,707 30/24000 2128 00:53:05
15,459,336 120/24000 2128 11:24:41
[…]
Table 7: Quantitative results on real-world downscaling. The evaluation is conducted on keyframes and interframes, where keyframes denote time-aligned LR–HR frame pairs, while interframes refer to intermediate time steps predicted by the neural network using two adjacent LR frames.
Frames Keyframe Interframes
LR Ours Bicubic Ours
Metric MAE MSE MAE MSE MAE MSE MAE MSE
0.0198 0.0052 0.0133 0.00251 0.0256 0.0057 0.0197 0.00295
To validate our method’s adaptation to a real-world flooding scenario, we generate a new dataset based on Hurricane Ike. The data are converted into video sequences of 205 frames at a resolution of $5803\times 5906$. We randomly split the data into training (165) and test (40) datasets. We use the training data to fine-tune our pretrained model for 100 epochs (approximately 2 hours). We then evaluate the model’s performance on the test dataset using the most recent model checkpoint. Figure 15 presents a visual comparison between the low-resolution (LR) inputs, our spatio-temporally downscaled results, and the high-resolution (HR) reference data under a real-world flooding scenario. Compared with the LR inputs, our method substantially enhances spatial details. It produces smoother, more coherent inundation structures that are visually consistent with the HR observations, particularly along the main flood fronts and large-scale flow patterns. The overall color gradients and shoreline geometries reconstructed by our model closely resemble those of the HR data, indicating effective recovery of fine-scale spatial variations during both moderate and extreme flooding stages. This observation is further supported by the residual visualizations (last row), in which the amplified errors remain confined mainly to narrow boundary regions and localized high-gradient areas, whereas the majority of flooded regions exhibit low residual magnitudes. In contrast, the LR inputs fail to capture these sharp transitions and localized structures. Although minor discrepancies persist in regions with complex topography or rapidly varying flood dynamics, the overall residual magnitude remains limited, suggesting that our method preserves both global consistency and local fidelity. These results demonstrate the strong capability of the proposed model to bridge the resolution gap between simulated inputs and HR flooding patterns, even in challenging real-world conditions. Quantitative comparison is shown in Table 7. We separately report the results of spatial downscaling (keyframes) and temporal downscaling (interframes). For spatial downscaling, paired LR and HR data are available, enabling direct comparison of the HR ground truth with the model predictions. For temporal downscaling, the model takes neighboring LR frames as input to estimate the intermediate frames via temporal interpolation. For comparison, we use temporal bicubic as the baseline method. As shown in the table, our method achieves improvements of approximately 0.002 in MSE and 0.006 in MAE relative to the baseline.